(9) Find the derivative of each of the following: (a) y = 3,5-1 – VI + V2 (d) y = sin-(In x) (8) y = In V272422+3 In 2x – 1 x Can u guys hlep me answer each of those questions please Thanks in advance. could you guys do them step by step and explain which rules u guys are used as you guys are doing the problems. please.
Answer #1

09 2 (a) 9=325.1 . 7 + T using formula 2² y=xn dy - sur 5.101 3x5.1X to I dy = nxn- dx differentiation of constant is zero ..(1) y = x + x dy - t* ln t + T x T-1 7 da Using dla) - aä lna dx o s d(xh) - nan da Rate Positive Please